OpenStudy (anonymous):

One square is placed inside another square so that their diagonals overlap. If the larger square has a side of 6 and the smaller square has a side of 4, how far is it from the vertex of the larger square to the corresponding vertex in the smaller square (marked in red)?

OpenStudy (anonymous):

\[\sqrt{2}\] \[2\sqrt{2}\] \[3\sqrt{2}\] \[4\sqrt{?}\]

OpenStudy (anonymous):

Using the pythagorean theorem to find the length of the diagonal in the 6x6 square and also in the 4x4 square, We can subtract those values to get 2x the red length, and we can just divide by two to find the red length

OpenStudy (anonymous):

so the answer will be \[\sqrt{2}\] right

OpenStudy (anonymous):

Thats what I got

OpenStudy (anonymous):

ok thANKS

OpenStudy (anonymous):

No problem